# Bioinspired Drone Rotors for Reduced Aeroacoustic Noise and Improved Efficiency

### v1

@@ 1 Introduction
The dominant source of noise from drones is aeroacoustic noise (Candeloro et al., 2022) and aeroacoustic noise primarily consists of tonal and broadband noise (Glegg and Devenport, 2017) components. Tonal noise is periodic and occurs mainly at the rotation frequency and its harmonics, and generally depends on blade thickness, shape, rotation speed and the advancing/retreating condition of the rotors. Reduction in tonal noise generally requires changing the shape of the blade or reducing rotational speeds but other more unconventional strategies have also been attempted (Boltezar et al., 1998, Kim, 2016). Broadband noise can be categorized as noise due to nonperiodic fluctuations and it is generated due to the flow separation, vortex shedding, blade-vortex interaction, and etc. All these flow structures affect the surface pressure of the blade and generate aeroacoustic noise. Computational studies on drone rotors show complex vortical structures (Misiorowski et al., 2019) and the interaction of the blade with these vortical structures is one of the major sources of the broadband noise. There have been several design modifications such as modifying the blade tip, the leading or trailing edge, blade planform shape and so on to try to mitigate the blade-vortex interaction noise (Yu, 2000). In some studies, bio-inspired designs have been employed to reduce the noise. The flight of the owls has been studied so that their salient features can be used to design silent UAVs. For instance, the wings of an owl contain features such as serrations and fringes to promote silent flight (Wagner et al., 2017) which was used by Noda et al. (2018) to develop a quieter drone propeller.
@@
The work in this paper is inspired by a study on the aeroacoustics of flapping flight in insects (Seo et al., 2019, 2021) where the authors compared the sound generated by flapping of mosquito and fruit fly wings. Despite comparable wing spans and weights, mosquitoes are known to generate a higher intensity sound from their flapping than fruit flies, and these studies examined the features of the wing morphology and kinematics that might explain this difference. The study employed direct numerical simulations of the flow coupled with an estimation of far-field sound via the Ffowcs William-Hawkings (FWH) acoustic analogy-based model and figure 1(a-b) shows the vortex structures for the two wings. Also shown in figure 1(c) is a comparison of the directivity pattern of the radiated aeroacoustic sound in terms of root-mean-square acoustic pressure normalized by the lift generated by the wing and the mosquito clearly generates a higher acoustic pressure per unit lift than the fruit fly. Indeed, as shown in the table 1, the specific acoustic power of the mosquito is about 3.7 times larger than a fruit fly.
@@
Thus, this comparative analysis suggests that an increase in the wetted area of the rotor blades accompanied by a reduction in the rotation speed could be an effective strategy for reducing the aeroacoustic sound from rotors for small-scale drones. Indeed, there is evidence that such a strategy could be useful even for larger-scale rotors. More than 70 years ago Vogeley (1949) demonstrated a reduction in aeroacoustic sound from an aircraft propeller using a similar idea. Figure 2 taken from this study shows the original two-blade nose propeller (figure 2(a)), which is replaced by a five-bladed propeller (figure 2(b)) where each propeller blade has a significantly larger wetted area and the radius of this propeller was also increased. Thus at a given airspeed of 55 mph, the modified blade propeller operated at a reduced rotation rate (794 RPM) compared to the unmodified propeller (2130 RPM) and this resulted in a reduction in the overall level of the aeroacoustic sound. Figure 2(c) shows a minimum reduction of about 15 dB in noise associated with this modification of the propellers and the authors mention that at highest airspeed and a given sound threshold of 65 dB, the unmodified plane has to be about 2000 ft away from the observer, while the modified plane could come as close as 200 ft. Interestingly, the modified aircraft also exhibited a higher mechanical efficiency, for instance, the mechanical efficiency with the modified propeller was 7.2% higher at an airspeed of 55 mph. These results suggest that similar design changes in drone rotors could lead to a reduction in acoustic emissions for a given lift while maintaining or even improving the power efficiency. This could result in quieter drones that are environmentally friendly and even increase the flight duration through improved efficiency.
@@ 2.2 Flow Solver
The rotor blade is represented by a zero-thickness plate comprised of unstructured triangular elements while the flow domain is discretized using a non-uniform Cartesian grid. The momentum equation uses Van-Kan’s fractional step method to decompose it into an advection-diffusion equation and a pressure Poisson equation. A second-order Adam-Bashforth scheme is used for the advection-diffusion equation, an implicit Crank-Nicholeson scheme is used for the viscous term, and biconjugate stabilized gradient descent is used for the pressure Poisson equation. The validation case studies for this code application to 2D and 3D cases can be found in Mittal et al. (2008) and the code has been used previously for complex cases involving rotating and flapping wing (Zheng et al., 2013).
@@ 2.3 Acoustics
To compute the noise generated by the rotors, we use the Ffowcs William-Hawkings (FWH, Ffowcs Williams et al. (1969)) equation and to make the computation less expensive, we specifically use the Brentner and Farassat integral formulation of the FW-H equation (Brentner and Farassat, 1998, 2003) given by
@@ 2.4 Scaling Analysis
The scaling in Eq. 6 indicates that for a constant span and a fixed lift force generated by the rotor, the rotor rotation speed has to satisfy the following relationship:
Combining the above with the expression in Eq. 10, we find that the acoustic intensity for a fixed lift would scale as:
This indicates that for the condition of fixed lift, the acoustic intensity reduces inversely with the blade area. Substituting the above scaling in the expression for aerodynamic power, we obtain the following:
@@ 3.1 Blade Pitch
We begin by simulating the rectangular blade at a pitch angle of for a range of values. Fig. 5(a) - 5(d) show the vortex structures for 4 of these cases and figure 5(e) shows the time variation of the lift over 4 cycles of rotation for four cases with ranging from 3,306 to 33,730. The domain size for these simulations were selected to be in the x, y and z directions respectively. In the z-direction, the uniform fine mesh region was confined to region (here, is the rotor chord length) placed such that below and region above the rotor center is resolved. The fine region in the plane of rotation (x-y plane) was box with the center of the box located at the center of the rotation and region surrounding the blade was uniform fine mesh and this was expanded very slowly away from the uniform region. The overall grid size ranged from 16 to 25 million grid points with more grid points used for larger area rotors. We conducted a grid refinement study, to ensure that the unsteady aerodynamic forces on the RR-10 blade were well converged on the current grid. We note that both the mean lift force as well as the fluctuations in the lift force increase with . The increase in mean lift will be discussed in detail in the following paragraph but the increase in fluctuation level is associated with the more intense unsteady vortex phenomena at higher and this is expected to increase the broadband component of the noise emitted from the blade.
@@
The sound spectrum for all four rotors is shown in figures 9(b) and 9(c) corresponding to polar locations of (6.5 , 0∘) and (6.5 , 90∘) respectively with the amplitude corresponding to RR-10 rotor subtracted from them to show the relative reduction or gain in the noise at each frequency. The larger area rotors rotate with lower frequency, and we observe the peak tonal noise being shifted to lower frequencies. We note that the rectangular blade which is operating at a higher has larger broadband noise and as the area of the blades is increased, the broadband noise is also reduced with a reduction in . This is an additional benefit of increased blade area.
@@ 3.3 Dual-Blade Rotors
In the previous sections, we have investigated a single rotating blade in order to simplify the analysis. While this verified the strategy of increasing blade area as a way to reduce noise and power consumption, single-blade rotors are seldom used and most drone rotors are dual-bladed. In this section, we conduct a limited study of this strategy for a dual-bladed rotor by comparing a dual bladed rectangular rotor RR-10 with the BI-20 rotor. The pitch angle chosen for this study is , which is more realistic for these rotors. The normalized rotation speed for these RR-10 and BI-20 rotors are 1.0 and 0.88 respectively to produce similar lift force, with the latter found via trial-and-error. The snapshot of the vortex structures for these rotors is shown in the figure 13(a) and 13(b). The RR-10 rotor shows a thin and long spiral shape tip-vortex being formed whereas BI-20 shows a more complex tip-vortex consisting of many vortex loops and it also occupies are larger region of the rotor wake.
@@
The lift and the lift coefficients are shown in the table 5 and we note that the difference between the mean lift produced by these rotors differs by about 3%. Similar to the single-blade rotor at a pitch angle of 20∘ case, we find a significant (nearly 65%) reduction in the lift coefficient of the BI-20 rotor compared to the RR-10 rotor. The directivity pattern shown in the figure 13(c) shows the dipole shape and we find that the BI-20 rotor generated lower noise in all the directions. The mean specific mechanical power and specific sound intensity (calculated at a distance of 6.5 from the rotation center) shown in table 5 indicates that the BI-20 rotor requires 15% less power and generates only 74% of the noise of the RR-10 rotor per unit lift.
@@ 4 Conclusions
Inspired by the aeroacoustics of insect flight we have examined a simple strategy to reduce the aeroacoustics noise from drone rotors - increased rotor blade area with fixed rotor span, that is accompanied by a concomitant decrease in rotation speed. The study employs direct numerical simulations of the rotor flow coupled with the prediction of aeroacoustic sound using the Ffowcs Williams-Hawkings acoustic model. Four different blades with shapes inspired by insect wings and with different total surface areas are employed. The initial set of simulations are for a single-bladed rotor and we have also examined the effect of blade pitch angle. The study design determines a rotation speed that generated a nearly equivalent mean lift for all the different rotors. The key aeroacoustic metric compared between the different blade designs is the aeroacoustic sound intensity per unit lift at a set distance from the center of the rotor and this is found to decrease with increasing blade area. It is also found that an additional benefit of this strategy is a reduction in the aerodynamic power per unit lift required to rotate the rotor. The study culminates with a comparison of a more realistic two-bladed rotor and reductions in aeroacoustic noise intensity and aerodynamic power are found for this case as well.

### v2

@@ 1 Introduction
The dominant source of noise from drones is aeroacoustic noise (Candeloro et al., 2022) that consists of tonal and broadband (Glegg and Devenport, 2017) components. Tonal noise is periodic and occurs mainly at the rotation frequency and its harmonics, and generally depends on blade thickness, shape, rotation speed and the advancing/retreating condition of the rotors. Reduction in tonal noise may require changing the shape of the blade or reducing rotational speed but other more unconventional strategies have also been attempted (Boltezar et al., 1998, Kim, 2016). Broadband noise can be categorized as noise due to non-periodic fluctuations and these are generated due to the blade interacting with turbulent flow structures. The broadband noise is particularly important at low tip Mach numbers (Hubbard, 1991) and this can be further sub-categorized as i) noise due to the flow separation, ii) noise due to the interaction of the shed vortices with other blades (blade-vortex interaction, BVI), iii) noise due to the interaction of boundary layer turbulence with trailing edge and leading edge. The incoming flow condition also affects the broadband noise (Petricelli et al., 2023). The tonal and the broadband noise together form the total loading noise which is the dominant noise components at low Mach flows (Glegg and Devenport, 2017). The aeroacoustics theories (Lighthill, 1952, Ffowcs Williams et al., 1969) showed that the source of the loading noise can be expressed by the surface pressure fluctuations. All the aforementioned flow structures affect the surface pressure of the blade and generate loading noise. Computational studies on drone rotors show complex vortical structures (Misiorowski et al., 2019) and the interaction of the blade with these vortical structures is one of the major sources of the broadband noise as mentioned above. There have been several design modifications such as modifying the blade tip, the leading or trailing edge, blade planform shape and so on to try to mitigate the blade-vortex interaction noise (Yu, 2000). In some studies, bio-inspired designs have been employed to reduce the noise. The flight of the owls has been studied so that their salient features can be used to design silent UAVs. For instance, the wings of an owl contain features such as serrations and fringes to promote silent flight (Wagner et al., 2017) which was used by Noda et al. (2018) to develop a quieter drone propeller.
@@
The work in this paper is inspired by a study on the aeroacoustics of flapping flight in insects (Seo et al., 2019, 2021) where the authors compared the sound generated by flapping of mosquito and fruit fly wings. Despite comparable wing spans and weights, mosquitoes are known to generate a higher intensity sound from their flapping than fruit flies, and these studies examined the features of the wing morphology and kinematics that might explain this difference. The study employed direct numerical simulations of the flow coupled with an estimation of far-field sound via the Ffowcs William-Hawkings (FWH) acoustic analogy-based model and figure 1(a-b) shows the vortex structures for the two wings. Also shown in figure 1c is a comparison of the directivity pattern of the radiated aeroacoustic sound in terms of root-mean-square acoustic pressure normalized by the lift generated by the wing and the mosquito clearly generates a higher acoustic pressure per unit lift than the fruit fly. Indeed, as shown in the table 1, the specific acoustic power of the mosquito is about 3.7 times larger than a fruit fly.
@@
Thus, this comparative analysis suggests that an increase in the wetted area of the rotor blades accompanied by a reduction in the rotation speed could be an effective strategy for reducing the aeroacoustic sound from rotors for small-scale drones. Indeed, there is evidence that such a strategy could be useful even for larger-scale rotors. More than 70 years ago Vogeley (1949) demonstrated a reduction in aeroacoustic sound from an aircraft propeller using a similar idea. Figure 2 taken from this study shows the original two-blade nose propeller (figure 2a), which is replaced by a five-bladed propeller (figure 2b) where each propeller blade has a significantly larger wetted area and the radius of this propeller was also increased. Thus at a given airspeed of 55 mph, the modified blade propeller operated at a reduced rotation rate (794 RPM) compared to the unmodified propeller (2130 RPM) and this resulted in a reduction in the overall level of the aeroacoustic sound. Figure 2c shows a minimum reduction of about 15 dB in noise associated with this modification of the propellers and the authors mention that at highest airspeed and a given sound threshold of 65 dB, the unmodified plane has to be about 2000 ft away from the observer, while the modified plane could come as close as 200 ft. Interestingly, the modified aircraft also exhibited a higher mechanical efficiency, for instance, the mechanical efficiency with the modified propeller was 7.2% higher at an airspeed of 55 mph. These results suggest that similar design changes in drone rotors could lead to a reduction in acoustic emissions for a given lift while maintaining or even improving the power efficiency. This could result in quieter drones that are environmentally friendly and even increase the flight duration through improved efficiency.
@@ 2.2 Flow Solver
The rotor blade is represented by a zero-thickness plate comprised of unstructured triangular elements while the flow domain is discretized using a non-uniform Cartesian grid. The momentum equation uses Van-Kan’s fractional step method to decompose it into an advection-diffusion equation and a pressure Poisson equation. A second-order Adam-Bashforth scheme is used for the advection-diffusion equation, an implicit Crank-Nicholeson scheme is used for the viscous term, and biconjugate stabilized gradient descent is used for the pressure Poisson equation.
@@ 2.3 Acoustics
To compute the noise generated by the rotors, we use the Ffowcs Williams-Hawkings (FWH, Ffowcs Williams et al. (1969)) equation and to make the computation less expensive, we specifically use the Brentner and Farassat integral formulation of the FW-H equation (Brentner and Farassat, 1998, 2003) given by
@@
2.4 Solver Validation
Seo et al. (2022) simulated a canonical case of circular cylinder at Re=100 with Mach number of 0.1 with this same solver and compared the results of FWH based acoustic solver for far field noise with direct acoustic numerical simulation (DaNS) showing an excellent match between the two. The solver has also been validated for the wing tones generated by the mosquito wing. Feugère et al. (2024) recorded the wing tones from a tethered mosquito. The wing motion was reconstructed from high-speed videos and the flow was simulated by ViCar3D and acoustics was predicted by using the FWH equation. The sound directivity pattern from the simulation and the experiment is reproduced from Feugère et al. (2024) in Fig. 5. The comparison is quite reasonable in terms of overall intensity and directivity thereby providing confidence in the simulation approach. The reader can find additional details and validation case studies for the ViCar3D solver in Mittal et al. (2008), Seo and Mittal (2011) and specific applications involving rotating and flapping wing can be found in Zheng et al. (2013), Mittal et al. (2025). The same flow-acoustic solver is used for the current rotor simulations. While we do not have experimental data for comparison and validation at the relatively low Reynolds that we are limited to in the current study, the above validation for similar flows combined with a grid verification study provides confidence in the accuracy of the results presented here.
@@ 2.5 Scaling Analysis
The scaling in Eq. 6 indicates that for a constant span and a fixed lift force generated by the rotor (which also implies a fixed disk loading), the rotor rotation speed has to satisfy the following relationship:
Combining the above with the expression in Eq. 10, we find that the acoustic intensity for a fixed lift (and disk loading) would scale as:
This indicates that for the condition of fixed value of lift (and also a fixed disk loading since the rotor disk area is the same for all the rotors here), the acoustic intensity reduces inversely with the blade area. Substituting the above scaling in the expression for aerodynamic power, we obtain the following:
@@ 3.1 Blade Pitch
We begin by simulating the rectangular blade at a pitch angle of for a range of values. Fig. 6(a) - 6(d) show the vortex structures for 4 of these cases and figure 6(e) shows the time variation of the lift over 4 cycles of rotation for four cases with ranging from 3,306 to 33,730. The domain size for these simulations were selected to be in the x, y and z directions respectively. In the z-direction, the uniform fine mesh region was confined to region (here, is the rotor chord length) placed such that below and region above the rotor center is resolved. The fine region in the plane of rotation (x-y plane) was box with the center of the box located at the center of the rotation and region surrounding the blade was uniform fine mesh and this was expanded very slowly away from the uniform region. The overall grid size ranged from 16 to 25 million grid points with more grid points used for larger area rotors. We conducted a grid refinement study, to ensure that the unsteady aerodynamic forces on the RR-10 blade were well converged on the current grid. We note that both the mean lift force as well as the fluctuations in the lift force increase with . The increase in mean lift will be discussed in detail in the following paragraph but the increase in fluctuation level is associated with the more intense unsteady vortex phenomena at higher and this is expected to increase the broadband component of the noise emitted from the blade.
@@
The sound spectrum for all four rotors is shown in figures 10(b) and 10(c) corresponding to polar locations of (6.5 , 0∘) and (6.5 , 90∘) respectively with the amplitude corresponding to RR-10 rotor subtracted from them to show the relative reduction or gain in the noise at each frequency. The larger area rotors rotate with lower frequency, and we observe the peak tonal noise being shifted to lower frequencies. We note that the rectangular blade which is operating at a higher has larger broadband noise and as the area of the blades is increased, the broadband noise is also reduced with a reduction in . This is an additional benefit of increased blade area.
@@ 3.3 Dual-Blade Rotors
In the previous sections, we have investigated a single rotating blade in order to simplify the analysis. While this verified the strategy of increasing blade area as a way to reduce noise and power consumption, single-blade rotors are seldom used and most drone rotors are dual-bladed. In this section, we conduct a limited study of this strategy for a dual-bladed rotor by comparing a dual bladed rectangular rotor RR-10 with the BI-20 rotor. The pitch angle chosen for this study is , which is more realistic for these rotors. The normalized rotation speed for these RR-10 and BI-20 rotors are 1.0 and 0.88 respectively to produce similar lift force, with the latter found via trial-and-error. This case was simulated for a total of four rotation cycles with the snapshot of the vortex structures for these rotors shown in the figures 14(a) and 14(b). The RR-10 rotor shows a thin and long spiral shape tip-vortex being formed whereas BI-20 shows a more complex tip-vortex consisting of many vortex loops and it also occupies are larger region of the rotor wake.
@@
The lift and the lift coefficients are shown in the table 5 (averaged between last two cycles) and we note that the difference between the mean lift produced by these rotors differs by about 3%. Similar to the single-blade rotor at a pitch angle of 20∘ case, we find a significant (nearly 65%) reduction in the lift coefficient of the BI-20 rotor compared to the RR-10 rotor. The directivity pattern shown in the figure 14(c) shows the dipole shape and we find that the BI-20 rotor generated lower noise in all the directions. The mean specific mechanical power and specific sound intensity (calculated at a distance of 6.5 from the rotation center) shown in table 5 indicates that the BI-20 rotor requires 15% less power and generates only 74% of the noise of the RR-10 rotor per unit lift.
@@ 4 Conclusions
Inspired by the aeroacoustics of flapping flight in insects, we have examined a simple strategy to reduce the aeroacoustics noise from drone rotors - increased rotor blade area with fixed rotor span, that is accompanied by a concomitant decrease in rotation speed. The study employs direct numerical simulations of the rotor flow coupled with the prediction of aeroacoustic sound using the Ffowcs Williams-Hawkings acoustic model. Four different blades with shapes inspired by insect wings and with different total surface areas are employed. The initial set of simulations are for a single-bladed rotor and we have also examined the effect of blade pitch angle. The study design determines a rotation speed that generated a nearly equivalent mean lift (and therefore an equivalent disk loading) for all the different rotors. The key aeroacoustic metric compared between the different blade designs is the aeroacoustic sound intensity per unit lift at a set distance from the center of the rotor and this is found to decrease with increasing blade area. It is also found that an additional benefit of this strategy is a reduction in the aerodynamic power per unit lift required to rotate the rotor. The study culminates with a comparison of a more realistic two-bladed rotor and reductions in aeroacoustic noise intensity and aerodynamic power are found for this case as well.